A square mirror has sides measuring 2 ft. less than the sides of a square painting. If the difference between their areas is 32 ft.^2, find the lengths of the sides of the mirror and the painting.
Answer by checkley77(12844)

X^2=(X+2)^2-32
X^2=X^2+4X+4-32
X^2-X^2-4X=-28
-4X=-28
X=-28/-4
X=7 FT. FOR THE SIDES OF THE SQUARE MIRROR.
7+2=9 FT. FOR THE SIDES OF THE SQUARE PAINTING.
PROOF:
7^2=9^2-32
49=81-32
49=49